OpenStudy (anonymous):

Please help solve 4 √486

OpenStudy (amistre64):

486^(1/4)

OpenStudy (amistre64):

4.695....

OpenStudy (amistre64):

or you want an exact solution ;)

OpenStudy (mathmagician):

\[3\sqrt[4]{6}\]

OpenStudy (anonymous):

486 | 2 243 | 3 081 | 3 027 | 3 486 = 2*(3^5) 009 | 3 003 | 3 001 | - \[\sqrt[4]{486} = \sqrt[4]{2*3^5} = 3\sqrt[4]{2*3} = 3\sqrt[4]{6}\]

OpenStudy (amistre64):

that was my next guess :)

OpenStudy (anonymous):

haha, just do it all.. lol